A 30.1 kg cart is rolling across a level surface at 12.5 m/s. Brakes are applied slowing the cart down to 10.0 m/s. Calculate work.
1
Expert's answer
2020-11-17T11:30:24-0500
"W=K_f-K_i\\W=0.5mv_f^2-0.5mv^2_i\\W=0.5(30.1)(12.5^2-10^2)=847\ J"